\begin{thebibliography}{aaaaaa}
%\bibitem[A]{accola} R. D. M. Accola, {\em{On Castelnuovo's inequality for algebraic curves. I}}, Transactions of the American Mathematical Society \textbf{251} (1979), 357-373.
%\bibitem[ACV]{twisted} D. Abramovich, A. Corti, and A. Vistoli, { \em{Twisted bundles and admissible covers}}, Communications in Algebra \textbf{31} (2003), 3547-3618.
\bibitem[A]{aprodu-higher} M. Aprodu, {\em{On the vanishing of higher syzygies of curves}}, Mathematische Zeitschrift \textbf{241} (2002), 1-15.
%\bibitem[Ap]{aprodu-remarks} M. Aprodu, {\em{Remarks on syzygies of $d$-gonal curves}}, Math. Res. Lett \textbf{12} (2005), 387-400.
%\bibitem[ACG]{ACGH1} E. Arbarello, M. Cornalba, P.A. Griffiths and J. Harris, {\em{Geometry of algebraic curves, Volume II}}, Springer \textbf{268}, 2011.
%\bibitem[ACGH]{ACG2} E. Arbarello, M. Cornalba, P.A. Griffiths and J. Harris, {\em{Geometry of algebraic curves, Volume I}}, Grundlehren der Mathematischen Wissenschaften \textbf{267}, Springer, Heidelberg (1985).

\bibitem[AF]{aprodu-farkas} M. Aprodu and G. Farkas, {\em{Green's Conjecture for smooth curves on arbitrary K3 surfaces}}, Compositio Math.\ \textbf{147} (2011), 839-851.
%\bibitem[AF]{aprodu-farkas-clay} M. Aprodu and G. Farkas, {\em{Koszul cohomology and applications to moduli}}, Grassmannians, moduli spaces and vector bundles, Clay Mathematics Proceedings \textbf{14} (2011), 25-50.

\bibitem[AFPRW]{AFPRW} M. Aprodu, G. Farkas, S. Papadima, C. Raicu and J. Weyman, {\em{Koszul modules and Green's conjecture}},  Inventiones Math.\ \textbf{218} (2019), 657-720.
%\bibitem[AK]{ara-kol} C. Araujo and J.  Koll\'{a}r, {\em{Rational Curves on Varieties}}, Higher Dimensional Varieties and Rational Points. Springer (2003).
%\bibitem[AH]{arbarello-harris} E. Arbarello and J. Harris, {\em{Canonical curves and quadrics of rank 4}}, Compositio Math. \textbf{43} (1981), 145-179.
%\bibitem[AM]{andreotti-mayer} A. Andreotti and A. L. Mayer, {\em{On period relations for abelian integrals on algebraic curves}}, Ann.\ Sc.\ Norm.\ Sup.\ Pisa (1967), 189-238.

%\bibitem[AN1]{aprodu-nagel-nonvanishing} M. Aprodu and  J. Nagel, {\em{Non-vanishing for Koszul cohomology of curves}}, Commentarii Math. Helv. \textbf{82} (2007), 617-628.

\bibitem[AN]{aprodu-nagel} M. Aprodu and  J. Nagel, {\em{Koszul cohomology and algebraic geometry}}, University Lecture Series \textbf{52}, American Mathematical Society, Providence, RI (2010).

%\bibitem[AM]{andreotti-mayer} A. Andreotti and A. L. Mayer, {\em{On period relations for abelian integrals on algebraic curves}}, Annali della Scuola Normale Superiore di Pisa \textbf{21} (1967),189-238.
%\bibitem[AS]{aprodu-sernesi} M. Aprodu and  E. Sernesi, {\em{Secant spaces and syzygies of special line bundles on curves}}, Algebra and Number Theory \textbf{9} (2015), 585-600.
%\bibitem[AS]{aprodu-sernesi-excess} M. Aprodu and  E. Sernesi, {\em{Excess dimension for secant loci in symmetric products of curves}}, Collect. Math. (2016). doi:10.1007/s13348-016-0166-2
%\bibitem[AV]{AV} M. Aprodu and C. Voisin, {\em{Green-Lazarsfeld's conjecture for generic curves of large gonality}}, C.R. Math. Acad. Sci. Paris \textbf{336} (2003), 335-339.
%\bibitem[AS]{arbarello-sernesi-petri} E. Arbarello and E. Sernesi, {\em{Petri's approach to the study of the ideal associated to a special divisor}}, Inventiones Math. \textbf{49} (1978), 99-119.
%\bibitem[AC1]{AC1} E. Arbarello, M. Cornalba, {\em{Su una congettura di Petri}}, Commentarii Math.\ Helvetici \textbf{56}, (1981),1--37.
%\bibitem[AC]{arbarello-cornalba} E. Arbarello, M. Cornalba, {\em{Footnotes to a paper of Beniamino Segre}}, Mathematische Annalen \textbf{256} (1981), 341-362.
%\bibitem[ACG]{ACG2} E. Arbarello, M. Cornalba, P.A. Griffiths, {\em{Geometry of algebraic curves, Volume II}}, Grundlehren der Mathematischen Wissenschaften \textbf{268}, Springer, Heidelberg (2011).

%\bibitem[A]{Aprodu} M. Aprodu, {\em{Green-Lazarsfeld gonality conjecture for a generic curve of odd genus}}, International Mathematics Research Notices \textbf{63} (2004), 3409-3416.
%\bibitem[AF]{aprodu-farkas} M. Aprodu and G. Farkas, {\em{Green's conjecture for curves on arbitrary K3 surfaces}}, Compositio Math.\ \textbf{147} (2011), 839-851.
%\bibitem[AF2]{AF-covers} M. Aprodu and G. Farkas, {\em{Green's Conjecture for general covers}}, Compact moduli spaces and vector bundles, Contemp. math. \textbf{564} (2012), 211-226.

%\cite[AH]{arbarello-harris} E. Arbarello and J. Harris, {\em{Canonical curves and quadrics of rank 4}}, Compositio Math. \textbf{43} (1981),
%145-179.
%\bibitem[AK]{alper-kresch} J. Alper and A. Kresch, {\em{Equivariant versal deformations of semistable curves}}, to appear in Michigan Math.\ J.
%\bibitem[AK1]{altman-kleiman-bertini} A. B. Altman and S. L. Kleiman, {\em{Bertini theorems for hypersurface sections containing a subscheme}}, Communications in Algebra \textbf{7}(1979), 775-790.
%\bibitem[AK2]{altman-kleiman} A. B. Altman and S. L. Kleiman, {\em{Compactifying the Picard scheme}}, Advances Math. \textbf{35} (1980), 50-112.



%\bibitem[AF]{clay} M. Aprodu and G. Farkas, {\em{Koszul cohomology and applications to moduli}}, Clay Mathematics Proceedings\ \textbf{14}, American Mathematical Society, Providence, RI (2011).

%\bibitem[AV]{AV} M. Aprodu and C. Voisin, {\em{Green-Lazarsfeld's conjecture for generic curves of large gonality}}, C.R. Math. Acad. Sci. Paris \textbf{336} (2003), 335-339.
%\bibitem[Bi]{birkar-topics} C.\ Birkar, {\em{Topics in algebraic geometry}}, arXiv:1104.5035.
\bibitem[Ba]{basili} B. Basili, {\em{Indice de Clifford des intersections compl\'etes de l'espace}}, Bull.\ Soc.\ math.\ \textbf{124}, (1996), 61-95.

\bibitem[Bi]{birkenhake}C.\ Birkenhake, {\em{Linear systems on projective spaces}}, Manuscripta Math. \textbf{88} (1995), 177-184.

\bibitem[BEGY]{erman-et-al} J. Bruce, D. Erman, S. Goldstein and J. Yang. {\em{Conjectures and computations about Veronese syzygies}}, Experimental Math.\ \textbf{29}, (2020), 398--413.

%\bibitem[BP]{bhosle-param} U. N. Bhosle and A. J. Parameswaran, {\em{Picard bundles and Brill–Noether loci in the compactified Jacobian of a nodal curve}}, Internat. Math. Res. Not. \textbf{15} (2014), 4241-4290.


%\bibitem[Ba]{babbage} D. W. Babbage, {\em{A note on the quadrics through a canonical curve}}, J. London Math. Soc. \textbf{1} (1939), 310-315.
%\bibitem[Be]{beauville-schottky} A. Beauville, {\em{Prym varieties and the Schottky problem}}, Invent. Math \textbf{41} (1977), 149--196.
%\bibitem[vB1]{bothmer-thesis} H-C Graf v Bothmer, {\em{Geometrische syzygien von kanonischen Kurven}}, Thesis, Universit\"at Bayreuth, 2000.
%\bibitem[vB2]{bothmer-preprint} H-C Graf v Bothmer, {\em{Geometric syzygies of canonical curves of even genus lying on a K3 surfaces}}, arXiv:math/0108078.

%\bibitem[vB3]{bothmer-JPAA} H-C Graf v Bothmer, {\em{Generic syzygy schemes}}, J.\ Pure and Applied Algebra \textbf{208} (2007), 867-876.
%\bibitem[vB4]{bothmer-Transactions} H-C Graf v Bothmer, {\em{Scrollar syzygies of general canonical curves with genus $\le8$}}, Transactions AMS \textbf{359} (2007), 465-488.
%\bibitem[B]{bourbaki-algebra} N. Bourbaki,  {\em{Algebra I: chapters 1-3}}, Volume 1, Springer, 1998.

%\bibitem[C1]{chen-rational} X. Chen, {\em{Rational curves on K3 surfaces}}, J. Alg. Geom. \textbf{8} (1999), 245-278.
%\bibitem[C]{chen-every-rational} X. Chen, {\em{A simple proof that rational curves on K3 are nodal}}, Math. Ann. \textbf{324} (2002), 71-104.
%\bibitem[C2]{chen-nodal} X. Chen, {\em{Nodal curves on K3 surfaces}}, arXiv:1611.07423.
%\bibitem[CFHR]{CFHR} F. Catanese, M. Franciosi, K. Hulek and M. Reid, {\em{Embeddings of curves and surfaces}}, Nagoya Math.\ J.\ \textbf{154} (1999), 185-220.
\bibitem[CP]{ciliberto-pareschi} C. Ciliberto and G. Pareschi, {\em{Pencils of minimal degree on curves on a K3 surface}}, J. Reine Ang. Math. \textbf{460} (1995), 15--36.
%\bibitem[CK]{ciliberto-knutsen} C. Ciliberto and A. L. Knutsen, {\em{On k-gonal loci in Severi varieties on general K3 surfaces and rational curves on hyperkähler manifolds}}, J.\  Math.\ Pures App.\ \textbf{101} (2014), 473-494.
%\bibitem[B]{beauville-stable} A. Beauville, {\em{Some stable vector bundles with reducible theta divisor}}, Manuscripta Math.\ \textbf{110} (2003), 343-349.
%\bibitem[B]{bopp} C. Bopp, {\em{Syzygies of $5$-gonal canonical curves}}, Documenta Mathematica \textbf{20} (2015), 1055-1069.
%\bibitem[B]{bujokas} G. Bujokas, {\em{The Hurwitz Space of Covers of an Elliptic Curve $E$ and the Severi Variety of Curves in $E \times \PP^1$}}, arXiv:1409.0927.
%\bibitem[BCGGM]{BCGGM} M. Bainbridge, D. Chen, Q. Gendron, S. Grushevsky and M. M\"oller, {\em{Compactification of Strata of Abelian Differentials}}, Duke Math.\ J.\, to appear.
%\bibitem[BE]{BE} D. Buchsbaum and D. Eisenbud, {\em{Generic free resolutions and a family of generically perfect ideals}}, Adv. Math \textbf{18} (1975), 245-301.
%\bibitem[BHT]{bogomolov-hassett-tschinkel} F. Bogomolov, B. Hassett and Y. Tschinkel, {\em{Constructing Rational Curves on K3 Surfaces}}, Duke Math.\ J.\ \textbf{157} (2011), 535-550.
%\bibitem[BS]{bopp-schreyer} C. Bopp and F.-O.Schreyer. {\em{A version of Green's conjecture in positive characteristic}}, arXiv:1803.10481.
%\bibitem[C1]{copp2} M. Coppens, {\em{On G.\ Martens' characterization of smooth plane curves}}, Bull.\ London Math.\ Soc.\ \textbf{20} (1988), 217-220.
%\bibitem[C2]{coppens} M. Coppens, {\em{The existence of k-gonal curves possessing exactly two linear systems $g^1_k$}}, Math. Annalen \textbf{307} (1997), 291-297.

\bibitem[CCK]{choi-kang-kwak} Y. Choi, P-L Kang and S. Kwak, {\em{Higher linear syzygies of inner projections}}, J. Algebra \textbf{305} (2006), 859-876.

%\bibitem[CE]{casnati-ekedahl} G. Casnati and T. Ekedahl, {\em{Covers of algebraic varieties I. A general structure theorem,
%covers of degree 3, 4 and Enriques surfaces}}, J. Algebraic Geometry \textbf{5} (1996), 439-460.
%\bibitem[CEFS]{CEFS} A. Chiodo, D. Eisenbud, G. Farkas and F.-O. Schreyer, {\em{Syzygies of torsion bundles and the geometry of the level $\ell$ modular variety over $\overline{\mathcal{M}}_g$}}, Invent.\ Math. \textbf{194} (2013), 73-118.

%\bibitem[CFVV]{CFVV} E. Colombo, G. Farkas, A. Verra and C. Voisin, {\em{Syzygies of Prym and paracanonical curves of genus 8}}, arXiv:1612.01026.

%\bibitem[C]{coppens} M. Coppens, {\em{One dimensional linear systems of type II on smooth curves}}, Ph.D. Thesis, Utrecht (1983).
%\bibitem[D]{deligne} P. Deligne, {\em{Le Lemme de Gabber}}, Ast\'erisque \textbf{127} (1985), 131-150.
%\bibitem[D]{debarre} O. Debarre, {\em{Sur le probleme de Torelli pour les vari\'et\'es de Prym}} American J.\ Math.\ \textbf{111} (1989), 111--134.
%\bibitem[DP]{deopurkar-patel} A. Deopurkar and A. Patel, {\em{The Picard rank conjecture for the Hurwitz spaces of degree up to five}}, Algebra and Number Theory \textbf{9} (2015), 459-492.
%\bibitem[DM]{donagi-morrison} R. Donagi, D. Morrison, {\em{Linear systems on K3 sections}}, Journal of Differential Geometry \textbf{29} (1989), 49-64.

\bibitem[Eh]{ehbauer} S. Ehbauer, {\em{Syzygies of points in projective space and applications}}, Zero-dimensional schemes. Proceedings of the international conference held in Ravelo, Italy. 1992.
%\bibitem[E]{eisenbud-orientation} D. Eisenbud, {\em{Green's conjecture: an orientation for algebraists}}, Free resolutions in commutative algebra and algebraic geometry (Sundance, UT, 1990) (1992), 51-78.

%\bibitem[E]{eisenbud-book} D. Eisenbud, {\em{Commutative Algebra: with a view toward algebraic geometry}}, Springer, 2013.
%\bibitem[Ei1]{eisenbud-determinantal} D. Eisenbud, {\em{Linear sections of determinantal varieties}}, American J.\ Math.\ \textbf{110} (1988), 541-575.

\bibitem[Ei]{eisenbud-syzygies} D. Eisenbud, {\em{The geometry of syzygies}}, Graduate Texts in Mathematics \textbf{229}, Springer-Verlag, New York, 2005.

\bibitem[EF]{ellia-franco} P.\ Ellia and D.\ Franco, {\em{Gonality, Clifford index and multisecants}}, J.\ Pure Appl.\ Alg.\ \textbf{158}, (2001), 25-39.
%\bibitem[EG]{eisenbud-goto} D. Eisenbud and S. Goto, {\em{Linear free resolutions and minimal multiplicity}}, J.\ Algebra \textbf{88} (1984), 89-133.
%\bibitem[EN]{EN} J. A. Eagon and D. G. Northcott, {\em{Ideals defined by matrices and a certain complex associated with them}}, Proc.\ R.\ Soc.\ Lond.\ A \textbf{269} (1962), 188-204.
\bibitem[EH]{eisenbud-harris-minimal} D. Eisenbud and J. Harris, {\em{On varieties of minimal degree (A centennial account)}}, Proc. Sympos. Pure Math. \textbf{46} (1987), 3--13.

%\bibitem[EP]{eisenbud-popescu} D. Eisenbud and S. Popescu, {\em{Syzygy ideals for determinantal ideals and the syzygetic Castelnuovo lemma}}, Commutative Algebra, Algebraic Geometry, and Computational Methods \textbf{1996}, 247-258.

%\bibitem[EH]{eisenbud-harris-limit} D. Eisenbud and J. Harris, {\em{Limit linear series: basic theory}}, Inventiones Math.\ \textbf{85} (1986), 337-371.
\bibitem[EL]{ein-lazarsfeld} L. Ein and R. Lazarsfeld, {\em{The gonality conjecture on syzygies of algebraic curves of
              large degree}}, Publ. Math. Inst. Hautes \'Etudes Sci. \textbf{122} (2015), 301-313.
 \bibitem[EL2]{ein-lazarsfeld-asymptotic} L. Ein and R. Lazarsfeld, {\em{Asymptotic Syzygies of Algebraic Varieties}}, Inventiones Math.\ \textbf{190} (2012), 603-646.
 %\bibitem[EG]{ed-gr} D. Edidin and W. Graham, {\em{Equivariant intersection theory}}, Inventiones Math. \textbf{131} (1998), 595-634.
%\bibitem[F1]{farkas-syzygies} G. Farkas, {\em{Syzygies of curves and the effective cone of $\mm _g$}}, Duke Math.\ J.\ \textbf{135} (2006), 53-98.
%\bibitem[F2]{farkas-koszul} G. Farkas, {\em{Koszul divisors on moduli spaces of curves}}, American J.\ Math.\ \textbf{31} (2009), 819-867.
%\bibitem[F]{farkas-progress} G. Farkas, {\em{Progress on syzygies of algebraic curves}}, Lecture Notes of the Unione Matematica Italiana \textbf{21} (Moduli of Curves, Guanajuato 2016), 107-138.
%\bibitem[FHL]{FHL} W. Fulton, J. Harris and R. Lazarsfeld, {\em{Excess linear series on an algebraic curve}}, Proc. American Math. Soc. \textbf{92} (1984), 320-322.

%\bibitem[FK1]{generic-secant} G. Farkas and M. Kemeny, {\em{The generic Green--Lazarsfeld Secant Conjecture}}, Inventiones Math. \textbf{203} (2016), 265-301.
%\bibitem[FK2]{high-level} G. Farkas and M. Kemeny, {\em{The Prym-Green Conjecture for torsion line bundles of high order}}, Duke Math. J.  \textbf{166} (2017) , 1103-1124.
\bibitem[FK]{lin-syz}  G. Farkas and M. Kemeny, {\em{Linear syzygies for curves of prescribed gonality}}. Advances in Math. \textbf{356} (2019), 106810.
%\bibitem[FK3]{res-odd} G. Farkas and M. Kemeny, {\em{The Resolution of Paracanonical Curves of Odd Genus}}, Geometry and Topology, \emph{to appear}.
%\bibitem[FL]{FaLu} G. Farkas and K. Ludwig, {\em{The Kodaira dimension of the moduli space of Prym varieties}}, J.\ Euro.\ Math. Society \textbf{12} (2010), 755-795.
%\bibitem[FuLa]{fulton-laz-connectedness} W. Fulton and R. Lazarsfeld, {\em{On the connectedness of degeneracy loci and special divisors}}, Acta Math.\ \textbf{146} (1981), 271-283.

%\bibitem[FMP]{farkas-mustata-popa} G. Farkas, M. Musta\c{t}\u{a} and M. Popa, {\em{Divisors on $\cM_{g, g+1}$ and the Minimal Resolution Conjecture for points on
%canonical curves}}, Annales Sci. de L'\'Ecole  Normale Sup\'erieure \textbf{36} (2003), 553-581.
%\bibitem[FP]{faber-pandharipande} C. Faber and R. Pandharipande, {\em{Relative maps and tautological classes}}, Journal of the European Mathematical Society \textbf{7} (2005), 13-49.
 
 %\bibitem[FaP]{faber-pand} C. Faber and R. Pandharipande, {\em{Relative maps and tautological classes.}} J.\ Euro.\ Math.\ Soc.\ \textbf{7} (2005),13-49.
%\bibitem[FP]{far-pand} G. Farkas and R. Pandharipande, {\em{The moduli space of twisted canonical divisors}}, to appear in J.\ Inst.\ Math.\ Jussieu.

%\bibitem[FR]{farkas-rimanyi} G. Farkas and R. Rim\'anyi, {\em{Quadric rank loci on moduli of curves and K3 surfaces}}, arXiv:1707.00756
%\bibitem[Fu]{fulton} W. Fulton, {\em{Intersection theory}}, Vol. 2, Springer, 2013.
%\bibitem[FT]{farkas-tarasca} G. Farkas and N. Tarasca, {\em{Du Val curves and the pointed Brill-Noether theorem}}, arXiv:1606.02725.
%\bibitem[GH]{griffiths-harris} P. Griffiths and J. Harris, {\em{On the variety of special linear systems on a general algebraic curve}}, Duke Math.\ J.\ \textbf{47}(1980), 233-272.

%\bibitem[Gi]{gieseker} D. Gieseker, {\em{Stable curves and special divisors: Petri's conjecture}}, Inventiones Math.\ \textbf{66}(1982), 251-275.
%\bibitem[Go]{gomez} T. G\'omez, {\em{Brill–Noether theory on singular curves and torsion-free sheaves on surfaces}}, Communications in Analysis and Geometry \textbf{9} (2001), 725-756.

%\bibitem[Gr1]{green-canonical} M. Green, {\em{The canonical ring of a variety of general type}}, Duke Math. J. \textbf{49} (1982), 1087-1113.
\bibitem[G1]{green-canonical} M. Green, {\em{The canonical ring of a variety of general type}}, Duke Math.\ J.\ \textbf{49} (1982), 1087-1113.
\bibitem[G2]{green-koszul} M. Green, {\em{Koszul cohomology and the cohomology of projective varieties}}, J.\  Differential Geo.\ \textbf{19} (1984), 125-171.
\bibitem[G3]{green-koszul-II} M. Green, {\em{Koszul cohomology and the cohomology of projective varieties. II. }}, Journal of Differential Geometry \textbf{20} (1984), 279-289.

% \bibitem[Gr3]{green-quadrics} M. Green, {\em{Quadrics of rank four in the ideal of a canonical curve}}, Inventiones Math.\ \textbf{75} (1984), 85-104.
% \bibitem[G3]{green-NL} M. Green, {\em{A new proof of the explicit Noether--Lefschetz theorem.}} J.\ Diff.\ Geo \textbf{27} (1988), 155-159.
%\bibitem[GL1]{green-lazarsfeld-projective} M. Green and R. Lazarsfeld, {\em{On the projective normality of complete linear series on an algebraic curve}}, Inventiones Math. \textbf{83} (1986), 73-90.
% \bibitem[GHS]{GHS} T. Graber, J. Harris and J. Starr, {\em{Families of Rationally Connected Varieties}}, J.\ American Math.\ Soc.\ \textbf{16} (2003), 57-67.
 %\bibitem[GL1]{green-lazarsfeld-special} M. Green and R. Lazarsfeld, {\em{Special divisors on curves on a K3 surface}}, Inventiones Math. \textbf{89} (1987), 357-370.
 
% \bibitem[GK]{gabai-kazez} D. Gabai and W. H. Kazez, {\em{The classification of maps of surfaces}}, Inventiones Math.\ \textbf{90} (1987), 219-242.
%\bibitem[HM]{ha-mu} J. Harris and D. Mumford, {\em{On the Kodaira dimension of the moduli space of curves}}, Inventiones Math \textbf{67} (1982), 23-86.
%\bibitem[H]{hartshorne-clifford} R.\ Hartshorne, {\em{Clifford Index of ACM Curves in $\PP^3$}} Milan J.\ Math.\ \textbf{70} (2002), 209-221.
\bibitem[H]{hartshorne} R. Hartshorne, {\em{Algebraic Geometry}}, Graduate Texts in Mathematics \textbf{52}, Springer-Verlag, New York, 1977.
%\bibitem[H2]{hartshorne-generalized} R. Hartshorne, {\em{Generalized divisors on Gorenstein curves and a theorem of Noether}}, J. Math. Kyoto Univ \textbf{26} (1986), 375-386.
\bibitem[HS]{hartshorne-schlesinger} R. Hartshorne and E. Schlesinger, {\em{Gonality of a general ACM curve in $\PP^3$}} Pacific J.\ math.\ \textbf{251} (2011), 269-313.
%\bibitem[Hu]{huybrechts-k3} D. Huybrechts, {\em{Lectures on K3 surfaces.}}, Cambridge University Press \textbf{158}, 2016.
%\bibitem[HL]{huybrechts-lehn} D. Huybrechts and M. Lehn, {\em{The geometry of moduli spaces of sheaves}}, Cambridge University Press, 2010.
%\bibitem[HLMP]{quadrics-rank3} K. Han, W. Lee, H. Moon and E. Park, {\em{Rank 3 Quadratic Generations of Veronese Embeddings}}, to appear in Compositio Math.

%\bibitem[HSV]{HSV} J. Herzog, A. Simis, and W. V. Vasconcelos, {\em{Approximation complexes of blowing-up rings.}} J.\ Algebra \textbf{74} (1982), 466-493.


%\bibitem[HM1]{ha-mo} J. Harris and I. Morrison, {\em{Moduli of Curves}}, Graduate Texts in Mathematics \textbf{187}, Springer-Verlag, New York, 1998.
%\bibitem[HM]{ha-mu} J. Harris and D. Mumford, {\em{On the Kodaira dimension of the moduli space of curves}}, Inventiones Math \textbf{67} (1982), 23-86.
%\bibitem[HR]{hirsch} A. Hirschowitz and S. Ramanan, {\em{New evidence for Green's Conjecture on syzygies of canonical curves}}, Annales Scientifiques de l'\'Ecole Normale Sup\'erieure \textbf{31} (1998), 145-152.

%\bibitem[JPW]{JPW} T. J\'ozefiak, P. Pragacz, J. Weyman, {\em{Resolutions of determinantal varieties and tensor complexes associated with symmetric and antisymmetric matrices}}, Ast\'erique \textbf{87} (1981), 109-189.

%\bibitem[J]{jongmans} F. Jongmans, {\em{Le probl\'eme des s\'eries speciales d'une courbe alg\'ebrique}}, Bull.\ Acad.\ R.\ Sci.\ Belgique \textbf{35} (1949), 1027-1041.
%\bibitem[K1]{kemeny-thesis} M. Kemeny, {\em{Stable Maps and Singular Curves on K3 surfaces}}, Thesis, Universit\"at Bonn (2015).
%\bibitem[Ka]{kass-survey} J. Kass, {\em{Singular curves and their compactified Jacobians}}, A Celebration of Algebraic Geometry, Clay Math Proc.\ \textbf{18}(2013), 391-428.

%\bibitem[K]{kovacs-rational} S. Kov\'acs, {\em{Rational Singularities}}, arXiv:1703.02269.


%\bibitem[K1]{betti-multiple} M. Kemeny, {\em{Betti Numbers of Curves and Multiple-Point Loci}}, arXiv:1804.09221.
%\bibitem[Ke1]{kemeny-moduli} M. Kemeny, {\em{The moduli of singular curves on K3 surfaces}}, J. Math. Pures App. \textbf{104} (2015), 882-920.

\bibitem[K1]{projecting} M. Kemeny, {\em{Projecting Syzygies of Curves}}, Algebraic Geometry \textbf{7} (2020), 561-580.
\bibitem[K2]{kemeny-voisin} M. Kemeny, {\em{Universal Secant Bundles and Syzygies of Canonical Curves}}, Inventiones Math.\ \textbf{223} (2021), 995-1026.
%\bibitem[KS]{koh-stillman} J. Koh, and M. Stillman, {\em{Linear syzygies and line bundles on an algebraic curve}}, J.\ Algebra \textbf{125} (1989), 120-132.

%\bibitem[L]{lascoux} A. Lascous, {\em{Syzygies des vari/'eti/'es d/'eterminales}}, Adv. Math.\ \textbf{30}(1978), 202-237. 
%\bibitem[L1]{lazarsfeld-BNP} R. Lazarsfeld, {\em{Brill-Noether-Petri without degenerations}}, J. Differential Geo.\ \textbf{23} (1986), 299-307.
%\bibitem[L2]{lazarsfeld-lectures} R. Lazarsfeld, {\em{Lectures on linear series}}, arXiv:alg-geom/9408011.



%\bibitem[O]{ogrady} K. O'Grady, {\em{Moduli of sheaves and the Chow group of K3 surfaces}}, J. Math. Pures App. \textbf{100}(2013), 701-718.





%\bibitem[K2]{kemeny-singular} M. Kemeny, {\em{The Moduli of Singular Curves on K3 Surfaces}}, J. Math. Pures. Appl. \textbf{104} (2015), 882-920.
%\bibitem[Ke]{Keem} C. Keem, {\em{On the variety of special linear systems on an algebraic curve}} Math.\ Annalen \textbf{288} (1990), 309--322.

%\bibitem[Kem]{kempf} G. Kempf, {\em{On the geometry of a theorem of Riemann}}, Annals of Mathematics (1973),178-185.

%\bibitem[Kl]{kleiman} S. Kleiman, {\em{Multiple-point formulae I: Iteration}}, Acta Mathematica \textbf{147} (1981), 13-49.

%\bibitem[KKZ]{KKZ} A. Kokotov, D. Korotkin and P. Zograf, {\em{Isomonodromic tau function on the space of admissible covers}}, Advances in Mathematics \textbf{227} (2011), 586-600.
%\bibitem[KM]{keel-mckernan} S. Keel and J. M$^c$Kernan, {\em{Contractible extremal rays on $\overline{\mathcal{M}}_{0,n}$}}, arXiv:alg-geom/9607009.
%\bibitem[L]{li-relative} J. Li, {\em{Stable morphisms to singular schemes and relative stable morphisms.}} J.\ Diff.\ Geo.\ \textbf{57} (2001), 509--578.
%\bibitem[LT]{li-tian} J. Li and G. Tian, {\em{Virtual moduli cycles and Gromov-Witten invariants}} J.\ American Math.\ Soc.\ \textbf{11} (1998), 119-174.


%\bibitem[G-Mar]{martens} G. Martens, {\em{On curves on K 3 surfaces}}, Algebraic Curves and Projective Geometry, Springer, Berlin (1989), 174--182. 
%\bibitem[H-Mar]{h-martens} H. Martens, {\em{On the varieties of special divisors on a curve}} J. Reine Ang. Math. \textbf{227} (1967), 111--120.
%\bibitem[May]{mayer} A. Mayer, {\em{Families of K3 surfaces}} Nagoya Math J. \textbf{48} (1972), 1-17.
%\bibitem[Mo]{morrison-large} D. Morrison, {\em{On K3 surfaces with large Picard number}} Invent. math. \textbf{75} (1984), 105-121.

%\bibitem[Mu]{mumford-prym} D. Mumford, {\em{Prym varieties I.}} Contributions to analysis. \textbf{1974}, 325-350.
%\bibitem[M]{mumford-quadratic} D. Mumford, {\em{Varieties defined by quadratic equations}}. Corso C.I.M.E. \textbf{1969} (Questions on
%algebraic varieties). Roma: Cremonese 1970.
%\bibitem[N]{noether} M. Noether {\em{Ueber die invariante Darstellung algebraischer Functionen}}, Math.\ Annalen \textbf{17} (1880), 263-284.
\bibitem[OP]{ottaviani-paoletti} G.\ Ottaviani and R.\ Paoletti, {\em{Syzygies of Veronese embeddings}}, Comp. Math. \textbf{125} (2001), 31-37.

%\bibitem[P]{patel-thesis} A. Patel, {\em{The Geometry of Hurwitz space}}, PhD thesis, Harvard University (2013).
%\bibitem[R]{rimanyi} R. Rim\'anyi, {\em{Multiple-point formulas?a new point of view}}, Pacific J. Math \textbf{202} (2002), 475--490.


%\bibitem[K2]{extremal-gonality} M. Kemeny, {\em{The extremal gonality conjecture for curves of arbitrary gonality}}, arXiv:1512.00212.
%\bibitem[Kn]{knutsen} A. Knutsen, {\em{On $k$-th order embeddings of K3 surfaces and Enriques surfaces}}, Manuscripta Math. \textbf{104} (2001), 211-237.
%\bibitem[M]{h-martens} H. Martens, {\em{On the varieties of special divisors on a curve}}, Journal f\"ur die reine und angewandte Mathematik \textbf{227} (1967), 111-120.
%\bibitem[Mu]{mumford} D. Mumford, {\em{Further comments on boundary points}}, AMS Summer School at
%Woods Hole (1964), (notes)
%\bibitem[Ma]{mayer} A.L. Mayer, {\em{Compactification of the variety of moduli of curves}}, Seminar, Institut for Advanced Study,
%Princeton (1969/70), (notes).
\bibitem[OP]{ott-pao} G. Ottaviani and R. Paoletti, {\em{Syzygies of Veronese embeddings}}. Comp.\ Math.\, \textbf{125} (2001), 31--37.
\bibitem[P]{park} J. Park, Jinhyung, {\em{Asymptotic vanishing of syzygies of algebraic varieties}}, Comm.\ American Math.\ Soc.\ \textbf{2} (2022), 133--148.
%\bibitem[P]{peeva} I. Peeva, {\em{Graded syzygies}}, Springer, 2010.
%\bibitem[R]{rego} C. J. Rego, {\em{The compactified Jacobian}}, Ann. Sci. \'Ec. Norm. Sup. \textbf{13} (1980), 211-223.
\bibitem[R]{raicu} C. Raicu, {\em{Representation stability for syzygies of line bundles on Segre--Veronese varieties}}, J.\ European Math.\ Soc.\ \textbf{18} (2016), 1201--1231.
%\bibitem[R]{rathmann} J. Rathmann, {\em{An effective bound for the gonality conjecture}}, arXiv:1604.06072.
%\bibitem[Ry]{rydh} D. Rydh {\em{The canonical embedding of an unramified morphism in an étale morphism}}, Math.\ Zeit.\ \textbf{268} (2011), 707-723.
%\bibitem[S1]{sagraloff} M. Sagraloff, {\em{Special linear series and syzygies of canonical curves of genus $9$}}, Thesis, Saarland University, 2005.
%\bibitem[S2]{sernesi-def} E. Sernesi, {\em{Deformations of algebraic schemes}}, Springer, 2007.
%\bibitem[SP]{stacks} The Stacks Project Authors, {\em{Stacks Project}}, \url{http://stacks.math.columbia.edu}, 2019.
%\bibitem[S]{segre} B. Segre, {\em{Sui moduli delle curve poligonali, e sopra un complemento al teorema di esistenza di Reimann}}, Math. Annalen \textbf{100} (1928), 537-551.
\bibitem[S]{schreyer} F.-O. Schreyer, {\em{Syzygies of canonical curves and special linear series}}, Math. Ann. \textbf{275} (1986), 105-137.
%\bibitem[SS]{sidman-smith} J. Sidman and G. Smith, {\em{Linear determinantal equations for all projective schemes}}, Algebra and Number Theory \textbf{5} (2011), 1041-1061.

%\bibitem[Sch2]{sch} F.-O. Schreyer, {\em{Green's conjecture for the general $p$-gonal curve of large genus}}, In: Algebraic curves and projective geometry, Springer Lecture Notes \textbf{1389} (1988), 254-260.
%\bibitem[Sch2]{schreyer-topics} F.-O. Schreyer, {\em{Some topics in computational algebraic geometry}}, In: Advances in algebra and geometry (Hyderabad, 2001), Hindustan Book Agency (2003), 263-278.
%\bibitem[SSW]{SSW} J. Schicho, F.-O. Schreyer and M. Weimann, {\em{Computational aspects of gonal maps and radical parametrization of curves}}, Appl.\ Algebra Engrg.\ Comm.\ Comput.\ \textbf{24} (2013), 313-341.
%\bibitem[Te]{montserrat} M. Teixidor i Bigas {\em{Green's conjecture for the generic $r$-gonal curve of genus $g\geq 3r-7$}} Duke Math.\ J.\ \textbf{111} (2002), 195-222.
%\bibitem[Ta]{tannenbaum} A. Tannenbaum, {\em{Families of curves with nodes on K3 surfaces}}, Math.\ Ann.\ \textbf{260} (1982), 239-253.
%\bibitem[Te]{Te} M. Teixidor, {\em{Syzygies using vector bundles}}, Transactions of the American Mathematical Society \textbf{359} (2007), 897-908.
%\bibitem[Ty]{Ty} I. Tyomkin, {\em{On Severi varieties on Hirzebruch surfaces}},  International Mathematical  Research Notices \textbf{23} (2007).
%\bibitem[vdGK]{vgeer-kouvidakis} G. van der Geer and A. Kouvidakis, { \em{The Hodge Bundle on Hurwitz Spaces}} Pure and Applied Mathematics Quarterly \textbf{7}, 1927-1308, (2011).
%\bibitem[V1]{voisin-NL} C. Voisin, {\em{Une pr\'ecision concernant le th\'eoreme de Noether}}, Math. Annalen \textbf{280} (1988), 605-611.
%\bibitem[V1]{voisin-deformation} C. Voisin, {\em{Deformation des Syzygies et Theorie de Brill?Noether}}, Proc.\ London Math.\ Soc.\ \textbf{3.3} (1993), 493-515.
\bibitem[V1]{V1} C. Voisin, {\em{Green's generic syzygy conjecture for curves of even genus lying on a $K3$ surface}}, J.\ European Math. Society \textbf{4} (2002), 363-404.
\bibitem[V2]{V2} C. Voisin, {\em{Green's canonical syzygy conjecture for generic curves of odd genus}}, Compositio Math.\ \textbf{141} (2005), 1163-1190.
%\bibitem[Wei]{weibel} C. Weibel, {\em{An introduction to homological algebra}}, Cambridge university press, 1995.
%\bibitem[Wei]{yi-wei} Y. Wei, {\em{Generic Green's Conjecture and Generic Geometric Syzygy Conjecture in Positive Characteristic}}, arXiv:2109.12187

%\bibitem[Wey1]{weyman-sym-ext} J. Weyman, {\em{Resolutions of the exterior and symmetric powers of a module}}, J.\ Algebra \textbf{58} (1979), 333-341.
%\bibitem[Wey2]{weyman-book} J. Weyman, {\em{Cohomology of vector bundles and syzygies}}, Cambridge University Press, 2003.
\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage{amsmath, amsthm, amssymb,amscd}
\usepackage{MnSymbol}
\usepackage{float}
\usepackage{graphicx}
\usepackage{xypic}
\usepackage[arrow, matrix, curve]{xy}
\usepackage{color}
\usepackage{enumitem}  
\usepackage{tikz-cd} 
\usepackage{url}

\usepackage{blindtext}

\usepackage[english]{babel}
\newcommand{\seq}{\subseteq}
\newcommand{\C}{\mathbb{C}}

\newtheorem{thm}{Theorem}[section]
\newtheorem*{thm-nl}{Theorem}
\newtheorem*{prop-nl}{Proposition}

\def\PP{{\mathbf P}}

\def\Pic0{{\rm Pic}^0(X)}

\newtheorem*{cor-nl}{Corollary}
\newtheorem{conjecture}[thm]{Conjecture}
\newtheorem*{conjecture-nl}{Conjecture}

\newtheorem*{quest-nl}{Question}
\newtheorem*{quests-nl}{Questions}

\newtheorem{prop}[thm]{Proposition}

\theoremstyle{remark}

\title{{Linear syzygies of projective space}}
\date{\today}

\author[M. Kemeny]{Michael Kemeny}

\address{University of Wisconsin-Madison, Department of Mathematics, 480 Lincoln Dr
\hfill \newline\texttt{}
 \indent WI 53706, USA} 
 \email{{\tt michael.kemeny@gmail.com}}

\begin{document}
\begin{abstract}
In this paper, we study the Ein--Lazarsfeld Conjecture of syzygies of Veronese varieties $\PP^n$ embedded by $\mathcal{O}_{\PP^n}(d)$. We show a vanishing statement which agrees with the conjecture up to highest and second-highest order for linear syzygies $q=1$.
\end{abstract}

\maketitle
\setcounter{section}{-1}
\section{Introduction}
In this paper, we study syzygies for $\PP^n$ embedded by $\mathcal{O}_{\PP^n}(d)$. The case $n=1$ is known, as then the resolution is an Eagon--Northcott complex \cite[Ch.\ 6]{eisenbud-syzygies}, so we take $n \geq 2$. The Ein--Lazarsfeld Conjecture of syzygies of a Veronese variety, \cite[Theorem B]{ein-lazarsfeld-asymptotic} says 
\begin{conjecture}[Ein--Lazarsfeld] \label{EL-conj}
Fix an index $1 \leq q \leq n$ and $d \geq n+1$, then
$$\mathrm{K}_{p,q}(\PP^n,\mathcal{O}_{\PP^n}(d)) \neq 0$$
if and only if 
$$\binom{d+q}{q}-\binom{d-1}{q}-q\leq p \leq \binom{d+n}{n}-\binom{d+n-q}{n-q}+\binom{n}{n-q}-q-1.$$
\end{conjecture}
The nonvanishing direction of the statement was shown in \cite{ein-lazarsfeld-asymptotic}, so we focus on the vanishing, which was previously considered by Ottaviani-Paoletti \cite{ott-pao}. Syzygies of Veronese varieties has been much studied in recent years, see for example \cite{park}, \cite{erman-et-al}, \cite{raicu}.\\

We first remark that existing results of Aprodu \cite{aprodu-higher}, Green \cite{green-koszul} and Birkenhake \cite{birkenhake} give Conjecture \ref{EL-conj} in the case $n=2$.
\begin{prop}[Aprodu, Green, Birkenhake]  \label{plane-thm}
Conjecture \ref{EL-conj} holds in the case $n=2$.
\end{prop}
We then show how the vanishings from this case generalize to statements about the Ein--Lazarsfeld Conjecture for $q=1$ and $q=n$, for all $n \geq 2$.
In the case $q=1$ of \emph{linear} syzygies, which have been studied in, for instance \cite{lin-syz} and \cite{ein-lazarsfeld}, the conjecture predicts the following vanishing:
\begin{conjecture} \label{EL-conj-linear}
Fix $n \geq 2$. For $d \geq n+1$, then
$$\mathrm{K}_{p,1}(\PP^n,\mathcal{O}_{\PP^n}(d)) = 0$$
if 
$$p \geq \binom{d+n-1}{n}+n-1.$$
\end{conjecture}
To obtain this from conjecture \ref{EL-conj}, we used Pascal's Triangle
$$\binom{d+n}{n}-\binom{d+n-1}{n-1}=\binom{d+n-1}{n}.$$

In this paper, we prove:
\begin{thm}  \label{main-thm}
Fix $n \geq 3$. For $d \geq 1$, then
$$\mathrm{K}_{p,1}(\PP^n,\mathcal{O}_{\PP^n}(d)) = 0$$
if 
$$p \geq \binom{d+n-1}{n}+\binom{d+n-2}{n-2}.$$
\end{thm}
Notice that the Theorem holds for any $d \geq 1$, whereas conjecture \ref{EL-conj-linear} requires $d \geq n+1$. Indeed, the Theorem for $d=1$ is very well known \cite[Proposition 1.5]{aprodu-nagel}, whereas for $n=2$, the result of conjecture \ref{EL-conj-linear} was already known for $d\geq 2$, \cite[Lemma 5.1]{aprodu-higher}. The nonvanishing arguments, on the other hand, in conjecture \ref{EL-conj}, require $d\geq n+1$. Also note that theorem  \ref{main-thm} requires $n\geq 3$ to even be stated.\\

 The bound in Theorem \ref{main-thm} is in general weaker than in the conjecture, but they only differ for terms of order less or equal to the third highest most order in $d$. For instance, for $n=3$, Theorem \ref{main-thm} states $\mathrm{K}_{p,1}(\PP^n,\mathcal{O}_{\PP^n}(d)) = 0$ for $p \geq \binom{d+2}{3}+d+n-1$ whereas the conjecture is $\mathrm{K}_{p,1}(\PP^n,\mathcal{O}_{\PP^n}(d)) = 0$ for $p \geq \binom{d+2}{3}+2$ (for $d\geq 4$).\\

 The proof of Theorem \ref{main-thm} uses the projection approach of Ehbauer \cite{ehbauer} and Aprodu \cite{aprodu-higher} to show Theorem \ref{main-thm} by induction on $d$, reduces to the trivial case $d=1$. The projection approach to syzygies has been used in \cite{projecting} and \cite{choi-kang-kwak}. As we explain in our proof, our result is optimal for our method of projection.\\

\textbf{Acknowledgements} I thank G.\ Farkas for a discussion on this topic. The author was supported by NSF grant DMS-2100782 and a Vilas Life Cycle award.
\smallskip

\section{Construction of Projection of Syzygies}
We begin with a construction of the projection and evaluation maps on syzygies, as in \cite[\S 2.2]{aprodu-nagel}, however we give it in more generality. We first construct and define the \emph{evaluation} map rather than the projection map, which may be constructed without an added assumption used in our construction of the projection map.
Let $X$ be a smooth, complex variety, let $N$ be a base-point free line bundle on $X$, and let $D \seq X$ be an effective divisor and assume further that the evaluation map
$$\mathrm{H}^0(X,N) \to \textrm{H}^0(N_{|_D})$$
is surjective. Let $L:=N(-D)$. Assume $L$ is also base point free on $X$. Consider the \emph{kernel bundles} $M_L$ and $M_N$ as in \cite[\S 2.1]{aprodu-nagel}, which fit into a commutative diagram
{\small{$$\begin{tikzcd}
& 0 \arrow[d] & 0 \arrow[d]  & 0 \arrow[d] \\
0 \arrow[r] & M_{L} \arrow[r] \arrow[d] & \mathrm{H}^0(X,L) \otimes \mathcal{O}_X \arrow[r] \arrow[d] &  L=N(-D) \arrow[r] \arrow[d] &0\\
0 \arrow[r] & M_N \arrow[r] \arrow[d] & \mathrm{H}^0(X,N) \otimes \mathcal{O}_X  \arrow[r] \arrow[d]  &  N \arrow[r] \arrow[d]& 0\\
0 \arrow[r] & \Gamma \arrow[r] \arrow[d] & \textrm{H}^0(N_{|_D})  \otimes \mathcal{O}_X  \arrow[r] \arrow[d]  &  N_{|_D}  \arrow[r] \arrow[d] & 0\\
& 0 & 0 & 0 &
\end{tikzcd},$$}}where $\Gamma$ is a vector bundle (since $D$ is a divisor) of rank $s:=\textrm{h}^0(N_{|_D})=\dim \textrm{H}^0(N_{|_D}).$ In particular we have the exact sequence of vector bundles
$$0 \to  M_{L} \to M_N \to \Gamma \to 0,$$
to which we can apply \cite[II. Exercise 5.16]{hartshorne}. For any integer $p>s$, we have an exact sequence of sheaves
$$0 \to F \to \bigwedge^p M_N \to \bigwedge^{p-s} M_L(-D)\to 0,$$
where $F$ is a coherent sheaf, using $\bigwedge^s \Gamma \simeq \mathcal{O}_X(-D).$ We have natural inclusions $M_L(-D) \seq M_L \seq M_N$, and so the surjective morphism $\bigwedge^p M_N \to \bigwedge^{p-s} M_L(-D)$ induces a morphism of sheaves $$ t_p\; : \; \bigwedge^p M_N \to \bigwedge^{p-s} M_N.$$ Up to scalar, the morphism $t_p$ can be described as such. We have a natural inclusion $\bigwedge^p M_N \seq \bigwedge^p \mathrm{H}^0(N) \otimes \mathcal{O}_X$, and there is a map
\begin{align*}
\alpha_p \; : \; \bigwedge^p \mathrm{H}^0(N)\otimes \mathcal{O}_X &\to \bigwedge^{p-s} \mathrm{H}^0(N) \otimes \mathcal{O}_X \\
v_1 \wedge \ldots \wedge v_p &\mapsto \sum_{\{i_1, \ldots, i_s\} \seq\{1, \ldots,p\}} (-1)^{i_1+\ldots+i_s}v_1 \wedge \ldots \wedge \hat{v}_{i_1} \wedge \ldots \wedge \hat{v}_{i_s}\wedge \ldots \wedge v_p \otimes \gamma(v_{i_1}\wedge \ldots \wedge_{i_s}),
\end{align*}
for $\gamma: \bigwedge^{s} \mathrm{H}^0(N)\otimes \mathcal{O}_X\to  \bigwedge^{s}\textrm{H}^0(N_{|_D}) \otimes \mathcal{O}_X\simeq \mathcal{O}_X$ (which is only uniquely defined up to scalar) and $\alpha_p (\bigwedge^p M_N) \seq \bigwedge^{p-s} M_N$. We have $t=\alpha_{|_{\bigwedge^p M_N}}$. After choosing the sign of the scalar defining $\gamma$ correctly, we have a commutative diagram 
{\small{$$\begin{tikzcd}
 \bigwedge^{p+1} \mathrm{H}^0(N)\; \; \; \arrow[d, "\beta_p"] \arrow[r, "\mathrm{H}^0(\alpha_{p+1})"] & \; \; \; \;  \bigwedge^{p+1-s} \mathrm{H}^0(N) \arrow[d, "\beta_{p-s}"]  & \\
  \mathrm{H}^0(\bigwedge^{p} M_{N}(N))\; \; \;\arrow[r, "\mathrm{H}^0(t_{p}\otimes \mathrm{id}_N)"]  & \;  \;  \; \; \mathrm{H}^0(\bigwedge^{p-s} M_{N}(N)) 
\end{tikzcd},$$}}
where the vertical maps are induced from taking global sections of the exact sequence
$$0 \to \bigwedge^i M_N \to  \bigwedge^i \mathrm{H}^0(N)\otimes \mathcal{O}_X \xrightarrow{b_i} \bigwedge^{i-1} M_N\otimes N \to 0,$$
for $i \in \{p+1, p+1-s \}$. Similarly to \cite[Remark 1.3]{aprodu-nagel}, the morphism of sheaves $b_i$ can be described as
$$v_1 \wedge \ldots \wedge v_i \mapsto \sum_j (-1)^j v_1 \wedge \ldots \wedge \hat{v}_j \wedge \ldots \wedge v_i \otimes \mathrm{ev}(v_j),$$
where $\mathrm{ev}$ is the evaluation morphism 
$$ \mathrm{ev} \; : \; \mathrm{H}^0(N) \otimes \mathcal{O}_X \to N.$$ From the commutative diagram, we get a morphism $$\mathrm{ev}_{p,D} \; : \; \mathrm{K}_{p,1}(X,N) \simeq \mathrm{Coker}(\beta_p) \to  \mathrm{K}_{p-s,1}(X,N) \simeq \mathrm{Coker}(\beta_p),$$
called the \emph{evaluation map}. When $s=1$, it coincides with the map defined in \cite[\S 2.2]{aprodu-nagel}. Up to scalar, it is the composition of $s$ evaluation maps
$$\mathrm{K}_{p-j,1}(X,N) \to \mathrm{K}_{p-j-1,1}(X,N), \; j \in \{1, \ldots, s\},$$
as in  \cite[\S 2.2]{aprodu-nagel}, with the morphism
$$\mathrm{H}^0(X,N) \to \mathrm{H}^0(X,N_{|_D})\simeq \C^s \to \C,$$
induced by the projection to the $s$ coordinates of $\C^s$.\\

In the case where, in addition $\mathrm{H}^1(\mathcal{O}_X)=0$, we can construct more. Then $\mathrm{K}_{p,1}(X,N) \simeq \mathrm{H}^1(\bigwedge^{p+1}M_N)$ and $\mathrm{H}^1(t_{p+1})$ induces, up to sign, the evaluation map
$$ \mathrm{K}_{p,1}(X,N)  \to  \mathrm{K}_{p-s,1}(X,N) .$$
Indeed we have a commutative diagram
{\small{$$\begin{tikzcd}
0 \arrow[r] &\bigwedge^{p+1} M_N \arrow[r] \arrow[d, "t_{p+1}"] &\bigwedge^{p+1} \mathrm{H}^0(N)\otimes \mathcal{O}_X \arrow[r, "b_{p+1}"] \arrow[d, "\alpha_{p+1}"]  &\bigwedge^{i-1} M_N\otimes N \arrow[r] \arrow[d, "t_p\otimes\mathrm{id}"] &0 \\
0 \arrow[r]  &\bigwedge^{p+1-s} M_N \arrow[r] &\bigwedge^{p+1-s} \mathrm{H}^0(N)\otimes \mathcal{O}_X \arrow[r, "b_{p+1-s}"]  &\bigwedge^{i-1} M_N\otimes N \arrow[r] &0 
\end{tikzcd}$$}}with exact rows, and then the claim follows from standard results about the functoriality of cohomology. As we saw above, the image of $t_{p+1}$ lands in $\bigwedge^{p+1-s}M_L(-D)$, so the evaluation map factors through a map
$$\textrm{pr}_{p,D} \; : \;  \\  \mathrm{K}_{p,1}(X,N)  \to  \mathrm{K}_{p-s,1}(X,L),$$
called the \emph{projection map} (this can be seen to coincide with the projection map of \cite{aprodu-nagel}, when $s=1$ and $M$ is the coordinate ring, using Ehbauer's theorem \cite{aprodu-higher}). Further, the image of $\textrm{pr}_{p,D}$ is contained in the image of $ \mathrm{H}^1(\bigwedge^{p+1-s}M_L(-D))) \to  \mathrm{H}^1(\bigwedge^{p+1-s}M_L) \simeq \mathrm{K}_{p-s,1}(X,L)$.

\section{Linear syzygies}
We first prove Proposition \ref{plane-thm} for the projective plane, i.e.\ $n=2$. For $q=2$ and $n=2$, the vanishing direction of the  Ein--Lazarsfeld Conjecture is equivalent to two statements, both known. The first is
$$\mathrm{K}_{p,2}(\PP^2,\mathcal{O}_{\PP^2}(d)) = 0$$
for $p > \binom{d+2}{2}-3$.\\

 By the Duality Theorem, \cite{green-koszul}, this statement is equivalent to 
$$\mathrm{K}_{\binom{d+2}{2}-3-p,1}(\PP^2,\mathcal{O}_{\PP^2}(-3);\mathcal{O}_{\PP^2}(d)) = 0$$
for $p > \binom{d+2}{2}-3$ which is trivial. The second statement is 
$$\mathrm{K}_{p,2}(\PP^2,\mathcal{O}_{\PP^2}(d)) = 0$$
for $p < 3d-2$, which is a result of Green and Birkenhake \cite{birkenhake}, with a statement generalized to a conjecture on condition $N_p$ for all $n$ by Ottaviani--Paoletti \cite{ottaviani-paoletti}.\\

So then, to prove the remainder of Conjecture \ref{EL-conj} in the case $n=2$, we must proving a vanishing of linear syzygies, namely we must prove the statement
$$\mathrm{K}_{p,1}(\PP^2,\mathcal{O}_{\PP^2}(d))=0,$$
for $d\geq 3$ and $p \geq \frac{1}{2}(d+2)(d+1)-d.$ 
\begin{proof}
As explained in the introduction, we need to show
$$\mathrm{K}_{p,1}(\PP^2,\mathcal{O}_{\PP^2}(d))=0,$$
for $d\geq 3$ and $p \geq \frac{1}{2}(d+2)(d+1)-d.$ This is \cite[Lemma 5.1]{aprodu-higher}.
\end{proof}

We now prove Theorem \ref{main-thm}.
\begin{proof}[Proof of Theorem \ref{main-thm}]
Fix an integer $n \geq 3$ and fix $d \geq 1$. We will apply the construction in the last section for $N=\mathcal{O}_{\PP^n}(d)$ and $D\in |\mathcal{O}_{\PP^n}(1)|$. Set $$s=h^0(\mathcal{O}_{\PP^n}(d))-h^0(\mathcal{O}_{\PP^n}(d-1))=h^0(\mathcal{O}_{\PP^{n-1}}(d))$$
by Pascal's triangle and consider $s$ general points $p_1, \ldots, p_s \in D$. Let $H \seq \PP(\mathrm{H}^0(\mathcal{O}_{\PP^n}(d)))$ be a hyperplane containing $p_1, \ldots, p_s$. 
As seen in the last section, we can view the evaluation map $$\mathrm{ev}_{p,D}: \mathrm{K}_{p,1}(\PP^n, \mathcal{O}_{\PP^n}(d)) \to \mathrm{K}_{p-s,1}(\PP^n, \mathcal{O}_{\PP^n}(d)) $$
as a composite of $s$ evaluation maps. By Aprodu \cite[\S 2, \S 3]{aprodu-higher}, $\mathrm{ev}_{p,D}$ is nonzero if we assume $ \mathrm{K}_{p,1}(\PP^n, \mathcal{O}_{\PP^n}(d))\neq 0$.  Since $\mathrm{H}^1(\mathcal{O}_{\PP^n})=0$, the image of $\mathrm{ev}_{p,D}$ is contained in the image of
$$\mathrm{H}^1(\bigwedge^{p+1-s}M_{\mathcal{O}_{\PP^n}(d-1)}(-1)) \to  \mathrm{K}_{p-s,1}(\PP^n, \mathcal{O}_{\PP^n}(d)) .$$
So $\mathrm{K}_{p,1}(\PP^n, \mathcal{O}_{\PP^n}(d)) \neq 0 \implies \mathrm{H}^1(\bigwedge^{p+1-s}M_{\mathcal{O}_{\PP^n}(d-1)}(-1)) \neq 0.$ But $\mathrm{H}^1(\mathcal{O}_{\PP^n}(-1)=0$, so $$ \mathrm{H}^1(\bigwedge^{p+1-s}M_{\mathcal{O}_{\PP^n}(d-1)}(-1)) \simeq \mathrm{K}_{p-s,1}(\PP^n; \mathcal{O}(-1), \mathcal{O}(d-1)),$$
by \cite[Remark 2.6]{aprodu-nagel} \footnote{Note that there is a typo in this book, the right hand side should state $\bigwedge^{p+1}W \otimes \mathrm{H}^1(X, F \otimes L^{q-1})$.}. But by the Vanishing Theorem \cite{green-koszul}, $\mathrm{K}_{p-s,1}(\PP^n; \mathcal{O}(-1), \mathcal{O}(d-1))=0$ for $p-s \geq \binom{d-2+n}{n}$. Note that $\mathrm{K}_{p-s,1}(\PP^n; \mathcal{O}(-1), \mathcal{O}(d-1))=\mathrm{K}_{p-s,0}(\PP^n; \mathcal{O}(d-2), \mathcal{O}(d-1))$, so this result is optimal by \cite[Remark 6.5]{ein-lazarsfeld-asymptotic}. But this gives
\begin{align*}
p &\geq \binom{d-2+n}{n}+s \\
&=\binom{d-2+n}{n}+\binom{d+n}{n}-\binom{d+n-1}{n}\\
&=\binom{d-2+n}{n}+\binom{d+n-1}{n-1}\\
&=\binom{d-1+n}{n}+\binom{d+n-1}{n-1}-\binom{d+n-2}{n-1}\\
&=\binom{d-1+n}{n}+\binom{d+n-2}{n-2}
\end{align*}
by Pascal's triangle, as required.
\end{proof}

We end this paper by considering Veronese syzygies in the other extreme, with $q=n$. In this case, one side of the vanishing statement is $$\mathrm{K}_{p,n}(\PP^n,\mathcal{O}_{\PP^n}(d)) = 0,$$ if 
$$p \geq \binom{d+n}{n}-n.$$ 
 By the Duality Theorem, \cite{green-koszul}, this statement is equivalent to 
$$\mathrm{K}_{\binom{d+n}{n}-n-1-p,1}(\PP^2,\mathcal{O}_{\PP^2}(-n-1);\mathcal{O}_{\PP^n}(d)) = 0$$
for $p \geq \binom{d+n}{n}-n$ which is trivial. So the vanishing in this case amounts to the following theorem of Ein--Lazrsfeld \cite[Remark 6.5]{ein-lazarsfeld-asymptotic} (in more generality), which we given a different proof of.
\begin{thm}[Ein--Lazarsfeld] \label{q=n-thm}
Fix $n \geq 2$. For $d \geq n+1$, then
$$\mathrm{K}_{p,n}(\PP^n,\mathcal{O}_{\PP^n}(d)) = 0$$
if 
$$p \leq \binom{d+n}{n}-\binom{d-1}{n}-n-1.$$
\end{thm}
\begin{proof}
Our proof of this theorem generalizes Ottaviani--Paoletti's proof of the statement in the special case $n=2$, \cite[Proposition 3.2]{ott-pao}. Note that this theorem requires $d\geq {n+1}$ to be stated so that $\binom{d-1}{n}$ has it's usual meaning.
The goal of this section is to prove Theorem \ref{q=n-thm}. 
Consider a smooth hypersurface $X \in |\mathcal{O}_{\PP^n}(d)|$. By the Lefschetz Theorem \cite{green-koszul}, it is equivalent to prove 
$$\mathrm{K}_{p,n}(X,\mathcal{O}_{X}(d)) = 0$$
for
$$p \leq \binom{d+n}{n}-\binom{d-1}{n}-n-1,$$
where $\mathcal{O}_{X}(d)$ denotes $\mathcal{O}_{\PP^n}(d)|_X$. Notice, by the adjunction formula, $$\omega_X \simeq \mathcal{O}_{X}(d-n-1).$$ By the duality theorem, whose assumptions clearly hold, we need
$$\mathrm{K}_{s-n-p,0}(X,\mathcal{O}_{X}(d-n-1); \mathcal{O}_{X}(d)) = 0$$
for $$s=h^0(\mathcal{O}_{X}(d))-1= \binom{d+n}{n}-1.$$
The claim now follows from the Vanishing Theorem  \cite{green-koszul}.
\end{proof}

\end{document}